# D-Brane Effective Lagrangian in Spacetimes with Boundaries

### v1

@@ 2 Bulk Lagrangian
1. When the $D_{p}$-brane is along the circle, the world-volume indices split as $a=(\tilde{a},y)$, and the spacetime coordinate field splits as $X^{\mu}(\sigma)=(X^{\tilde{\mu}}(\sigma),y)$. The resulting reduction of the Lagrangian is denoted as $\mathbf{L}_{p}^{w}$.
2. When the $D_{p}$-brane is orthogonal to the circle, the world-volume indices do not split, i.e., $a=\tilde{a}$, and the spacetime coordinate field splits as $X^{\mu}(\sigma)=(X^{\tilde{\mu}}(\sigma),X^{y}(\sigma))$. The corresponding reduction of the Lagrangian is denoted as $\mathbf{L}_{p}^{t}$.
These two reductions are not identical. However, the transformation of $\mathbb{L}_{p}^{w}$ under the T-duality transformations
@@
which is called $\mathbb{L}_{p-1}^{wT}$, should be the same as $\mathbb{L}_{p-1}^{t}$. One can easily find that for the following reduction for the metric:
@@
where $g_{\tilde{\mu}\tilde{\nu}}$ is the constant base space metric and 1 is the radius of the circle, the reductions $\mathbf{L}_{p}^{w}$ and $\mathbf{L}_{p-1}^{t}$ for the Lagrangian in (4) are
$\mathbb{L}_{p}^{(0)w}$ $=$ $-T_{p}\sqrt{-\det({\tilde{g}}_{{\tilde{a}}{\tilde{b}}}+F_{{\tilde{a}}{\tilde{b}}}+\partial_{\tilde{a}}A_{y}\partial_{\tilde{b}}A_{y})}\,,$ (9)
$\mathbb{L}_{p-1}^{(0)t}$ $=$ $-T_{p-1}\sqrt{-\det({\tilde{g}}_{{\tilde{a}}{\tilde{b}}}+F_{{\tilde{a}}{\tilde{b}}}+\partial_{\tilde{a}}S\partial_{\tilde{b}}S)}\,,$
@@ 2.2 T-Duality Constraint on the Basis
where $\tilde{\nabla}$ is covariant derivative construncted from the base space pull-back metric ${\tilde{g}}_{{\tilde{a}}{\tilde{b}}}$. In particular, the only non-zero component of the second fundamental form is $\Omega_{{\tilde{a}}{\tilde{b}}}{}^{{\tilde{\mu}}}=\tilde{\nabla}_{\tilde{a}}\partial_{\tilde{b}}{}X^{\tilde{\mu}}\equiv\tilde{\Omega}_{{\tilde{a}}{\tilde{b}}}{}^{{\tilde{\mu}}}$. In this way, one can easily calculate $\mathbb{L}_{p}^{(1)w}$. Then using the transformation (7), one can calculate $\mathbb{L}_{p}^{(1)wT}$.
In the second case where $D_{p}$-brane is orthogonal to the circle, there is no world-volume index $y$. So the $F$ and pull-back metric has only world-volume indices. The pull-back metric is
@@
However, there is no such relation for $\tilde{\Omega}_{{\tilde{a}}{\tilde{b}}}{}^{y}$, i.e., $\tilde{\Omega}_{{\tilde{a}}{\tilde{b}}}{}^{y}\partial_{\tilde{c}}S\neq 0$. Using the above steps, one can calculate $\mathbb{L}_{p}^{(1)t}$.
If one imposes the constraint that $\mathbb{L}_{p}^{(1)wT}$ is the same as $\mathbb{L}_{p}^{(1)t}$, then one would find the wrong result that all coupling constants in (19) are zero. This indicates that the T-duality transformation (7) must recive higher-derivative corrections. Since we did not use $X^{\mu}$-field redefinition to find the independent couplings in (19), the T-duality also hase no correction to the last line in (7). For the other fields, we consider the following corrections:
@@
${A_{\tilde{a}}}$$\rightarrow$ ${A_{\tilde{a}}}+\sqrt{\alpha^{\prime}}\delta A_{\tilde{a}},$
${X^{\tilde{\mu}}}$$\rightarrow$ ${X^{\tilde{\mu}}}\,.$ (25)
@@
On the other hand, the transformation of $\mathbb{L}_{p}^{(0)w}$ under the above T-duality transformation produces $\mathbb{L}_{p-1}^{(0)t}$ on the second line of (9) and the following couplings at order $\alpha^{\prime}$:
$\delta\mathbb{L}_{p-1}^{(0)t}=-\frac{T_{p-1}\alpha^{\prime}}{2}\sqrt{-\det({\tilde{g}}_{{\tilde{a}}{\tilde{b}}}+F_{{\tilde{a}}{\tilde{b}}}+\partial_{\tilde{a}}S\partial_{\tilde{b}}S)}{\rm Tr}\Big{[}({\tilde{g}}+F+\partial S\partial S)^{(-1)}(\delta F+\delta(\partial S\partial S)\Big{]}\,,$ (28)
@@ 3 Boundary Lagrangian
The world-volume reduction of the Lagrangian in (33) then becomes
$\partial\mathbb{L}_{p}^{w}=-T_{p}\,e^{-\bar{\phi}+\varphi/4}\sqrt{-\det\Big{[}\partial_{\tilde{a}}X^{\tilde{\mu}}\partial_{\tilde{b}}X^{\tilde{\nu}}(\bar{g}_{{\tilde{\mu}}{\tilde{\nu}}}+\bar{b}_{{\tilde{\mu}}{\tilde{\nu}}}+2\sqrt{\alpha^{\prime}}\bar{K}^{\prime}_{{\tilde{\mu}}{\tilde{\nu}}})+F_{{\tilde{a}}{\tilde{b}}}+e^{-\varphi}\partial_{\tilde{a}}A_{y}\partial_{\tilde{b}}A_{y}\Big{]}}\,.$ (37)
And the transverse reduction of the Lagrangian in (33) becomes
$\partial\mathbb{L}_{p-1}^{t}=-T_{p-1}\,e^{-\bar{\phi}-\varphi/4}\sqrt{-\det\Big{[}\partial_{\tilde{a}}X^{\tilde{\mu}}\partial_{\tilde{b}}X^{\tilde{\nu}}(\bar{g}_{{\tilde{\mu}}{\tilde{\nu}}}+\bar{b}_{{\tilde{\mu}}{\tilde{\nu}}}+2\sqrt{\alpha^{\prime}}\bar{K}^{\prime}_{{\tilde{\mu}}{\tilde{\nu}}})+F_{{\tilde{a}}{\tilde{b}}}+e^{\varphi}\partial_{\tilde{a}}S\partial_{\tilde{b}}S\Big{]}}\,.$ (38)
The transformation of $\partial\mathbb{L}_{p}^{w}$ under T-duality (7) becomes identical to $\partial\mathbb{L}_{p-1}^{t}$.
@@ 4 Discussion
where $K^{\prime}_{\mu\nu}\equiv K_{\mu\nu}-2n_{\mu}\partial_{\nu}\Phi$. The corrections $O_{bulk}(\alpha^{\prime})$ for the bulk and boundary are the same. In the pull-back of the spacetime fields in the bulk, $X^{\mu}=X^{\mu}(\sigma)$, and in the boundary, $X^{\mu}=X^{\mu}(\tau)$. The bulk world-volume coordinates are $\sigma^{0},\cdots,\sigma^{p}$, and the boundary world-volume coordinates are $\tau^{0},\cdots,\tau^{p}$. Each term in the $\alpha^{\prime}$-corrections $O_{n}(\alpha^{\prime})$ should involve the unit vector $n_{\mu}$.
@@
We have found the T-duality invariant couplings (32) at order $\alpha^{\prime}$ up to the 8th order of the dimensionless field strength $F$ for the particular basis (19) in which the covariant derivatives are constructed from the pull-back metric ${\tilde{G}}_{ab}$ and indices are also contracted with inverse of this metric. In principle, one may extend this ${\tilde{G}}$-covariant basis to include all orders of $F$ and impose T-duality to find the corresponding coupling constants. This raises the question of whether it is possible to find a particular scheme in which all orders of $F$ can be written in a closed-form expression. A possible way to find such a scheme may be the following: The Lagrangian (32) has an overall DBI factor, which is the square root of the following metric:
@@
On the other hand, if one chooses another scheme for the basis (19), it is possible to change the coefficient of the last term in the second line to be 2, and remove the first term in the fifth line of (32). Then all terms in (32) with the structure $\Omega_{ab}{}^{\mu}\Omega_{cd\mu}$ whose world-volume indices are contracted with an even number of $F$, i.e., all terms in the first and second lines and the first term in the third line, can be written as $-{\tilde{{\cal G}}}^{ac}{\tilde{{\cal G}}}^{bd}\Omega_{ab}{}^{\mu}\Omega_{cd\mu}$, where ${\tilde{{\cal G}}}^{ab}$ is the inverse of the above metric. Therefore, one may expect that the world-volume covariant derivative constructed from the above metric, i.e., $\nabla_{{\tilde{{\cal G}}}}$, may be the appropriate covariant derivative to consider for constructing the second fundamental form and constructing the independent basis in which all indices are contracted with the inverse of the above metric. Then T-duality can fix the coupling constants of such a ${\tilde{{\cal G}}}$-covariant basis. It would be interesting to perform the appropriate calculation to determine the coupling constants of the ${\tilde{{\cal G}}}$-covariant basis. It may be possible that the infinite number of couplings at order $\alpha^{\prime}$ can be expressed as a few couplings constructed from the ${\tilde{{\cal G}}}$-metric.

### v2

@@ 2 Bulk Lagrangian
1. When the $D_{p}$-brane is along the circle, the world-volume indices split as $a=(\tilde{a},y)$, and the spacetime coordinate field splits as $X^{\mu}(\sigma)=(X^{\tilde{\mu}}(\sigma),y)$. The resulting reduction of the Lagrangian is denoted as ${L}_{p}^{w}$.
2. When the $D_{p-1}$-brane is orthogonal to the circle, the world-volume indices do not split, i.e., $a=\tilde{a}$, and the spacetime coordinate field splits as $X^{\mu}(\sigma)=(X^{\tilde{\mu}}(\sigma),X^{y}(\sigma))$. The corresponding reduction of the Lagrangian is denoted as ${L}_{p-1}^{t}$.
These two reductions are not identical. However, the transformation of ${L}_{p}^{w}$ under the T-duality transformations
@@
which is called ${L}_{p-1}^{wT}$, should be the same as ${L}_{p-1}^{t}$. One can easily find that for the following reduction for the metric:
@@
where $g_{\tilde{\mu}\tilde{\nu}}$ is the constant base space metric and 1 is the radius of the circle, the reductions ${L}_{p}^{w}$ and ${L}_{p-1}^{t}$ for the Lagrangian in (4) are
${L}_{p}^{(0)w}$ $=$ $-T_{p}\sqrt{-\det({\tilde{g}}_{{\tilde{a}}{\tilde{b}}}+F_{{\tilde{a}}{\tilde{b}}}+\partial_{\tilde{a}}A_{y}\partial_{\tilde{b}}A_{y})}\,,$ (9)
${L}_{p-1}^{(0)t}$ $=$ $-T_{p-1}\sqrt{-\det({\tilde{g}}_{{\tilde{a}}{\tilde{b}}}+F_{{\tilde{a}}{\tilde{b}}}+\partial_{\tilde{a}}S\partial_{\tilde{b}}S)}\,,$
@@ 2.2 T-Duality Constraint on the Basis
where $\tilde{\nabla}$ is covariant derivative construncted from the base space pull-back metric ${\tilde{g}}_{{\tilde{a}}{\tilde{b}}}$. In particular, the only non-zero component of the second fundamental form is $\Omega_{{\tilde{a}}{\tilde{b}}}{}^{{\tilde{\mu}}}=\tilde{\nabla}_{\tilde{a}}\partial_{\tilde{b}}{}X^{\tilde{\mu}}\equiv\tilde{\Omega}_{{\tilde{a}}{\tilde{b}}}{}^{{\tilde{\mu}}}$. In this way, one can easily calculate ${L}_{p}^{(1)w}$. Then using the transformation (7), one can calculate ${L}_{p-1}^{(1)wT}$.
In the second case where $D_{p-1}$-brane is orthogonal to the circle, there is no world-volume index $y$. So the $F$ and pull-back metric has only world-volume indices. The pull-back metric is
@@
However, there is no such relation for $\tilde{\Omega}_{{\tilde{a}}{\tilde{b}}}{}^{y}$, i.e., $\tilde{\Omega}_{{\tilde{a}}{\tilde{b}}}{}^{y}\partial_{\tilde{c}}S\neq 0$. Using the above steps, one can calculate ${L}_{p-1}^{(1)t}$.
If one imposes the constraint that ${L}_{p-1}^{(1)wT}$ is the same as ${L}_{p-1}^{(1)t}$, then one would find the wrong result that all coupling constants in (19) are zero. This indicates that the T-duality transformation (7) must recive higher-derivative corrections. Since we did not use $X^{\mu}$-field redefinition to find the independent couplings in (19), the T-duality also has no correction to the last line in (7). For the other fields, we consider the following corrections:
@@
${A_{\tilde{a}}}$$\rightarrow$ ${A_{\tilde{a}}}+\sqrt{\alpha^{\prime}}\delta A_{\tilde{a}}\,.$ (25)
@@
On the other hand, the transformation of ${L}_{p}^{(0)w}$ under the above T-duality transformation produces ${L}_{p-1}^{(0)wT}={L}_{p-1}^{(0)t}$ on the second line of (9) and the following couplings at order $\alpha^{\prime}$:
$\delta{L}_{p-1}^{(0)wT}=-\frac{T_{p-1}\alpha^{\prime}}{2}\sqrt{-\det({\tilde{g}}_{{\tilde{a}}{\tilde{b}}}+F_{{\tilde{a}}{\tilde{b}}}+\partial_{\tilde{a}}S\partial_{\tilde{b}}S)}{\rm Tr}\Big{[}({\tilde{g}}+F+\partial S\partial S)^{(-1)}(\delta F+\delta(\partial S\partial S)\Big{]}\,,$ (28)
@@ 3 Boundary Lagrangian
The world-volume reduction of the $D_{p}$-brane Lagrangian in (33) then becomes
$\partial{L}_{p}^{w}=-T_{p}\,e^{-\bar{\phi}+\varphi/4}\sqrt{-\det\Big{[}\partial_{\tilde{a}}X^{\tilde{\mu}}\partial_{\tilde{b}}X^{\tilde{\nu}}(\bar{g}_{{\tilde{\mu}}{\tilde{\nu}}}+\bar{b}_{{\tilde{\mu}}{\tilde{\nu}}}+2\sqrt{\alpha^{\prime}}\bar{K}^{\prime}_{{\tilde{\mu}}{\tilde{\nu}}})+F_{{\tilde{a}}{\tilde{b}}}+e^{-\varphi}\partial_{\tilde{a}}A_{y}\partial_{\tilde{b}}A_{y}\Big{]}}\,.$ (37)
And the transverse reduction of the $D_{p-1}$-brane Lagrangian in (33) becomes
$\partial{L}_{p-1}^{t}=-T_{p-1}\,e^{-\bar{\phi}-\varphi/4}\sqrt{-\det\Big{[}\partial_{\tilde{a}}X^{\tilde{\mu}}\partial_{\tilde{b}}X^{\tilde{\nu}}(\bar{g}_{{\tilde{\mu}}{\tilde{\nu}}}+\bar{b}_{{\tilde{\mu}}{\tilde{\nu}}}+2\sqrt{\alpha^{\prime}}\bar{K}^{\prime}_{{\tilde{\mu}}{\tilde{\nu}}})+F_{{\tilde{a}}{\tilde{b}}}+e^{\varphi}\partial_{\tilde{a}}S\partial_{\tilde{b}}S\Big{]}}\,.$ (38)
The transformation of $\partial{L}_{p}^{w}$ under T-duality (7) becomes identical to $\partial{L}_{p-1}^{t}$.
@@ 4 Discussion
where $K^{\prime}_{\mu\nu}\equiv K_{\mu\nu}-2n_{\mu}\partial_{\nu}\Phi$. The corrections $O_{bulk}(\alpha^{\prime})$ for the bulk and boundary are the same. In the pull-back of the spacetime fields in the bulk, $X^{\mu}=X^{\mu}(\sigma)$, and in the boundary, $X^{\mu}=X^{\mu}(\tau)$. The bulk world-volume coordinates are $\sigma^{0},\cdots,\sigma^{p}$, and the boundary world-volume coordinates are $\tau^{0},\cdots,\tau^{p-1}$. Each term in the $\alpha^{\prime}$-corrections $O_{n}(\alpha^{\prime})$ should involve the unit vector $n_{\mu}$.
@@
We have found the T-duality invariant couplings (32) at order $\alpha^{\prime}$ up to the 8th order of the dimensionless field strength $F$ for the particular basis (19) in which the covariant derivatives are constructed from the pull-back metric ${\tilde{G}}_{ab}$ and indices are also contracted with inverse of this metric. In principle, one may extend this ${\tilde{G}}$-covariant basis to include all orders of $F$ and impose T-duality to find the corresponding coupling constants. This raises the question of whether it is possible to find a particular scheme in which all orders of $F$ can be written in a closed-form expression. One approach to finding such a scheme might involve considering the Lagrangian (32), which features an overall DBI factor. This factor is given by the square root of the inverse of the following matrix:
$\left(\frac{1}{{\tilde{G}}+F}\right)^{ab}$ $=$ ${\tilde{{\cal G}}}^{ab}-\Theta^{ab}\,,$ (44)
where the symmetric and antisymmetric parts are, respectively,
${\tilde{{\cal G}}}^{ab}=\left(\frac{1}{{\tilde{G}}+F}{\tilde{G}}\frac{1}{{\tilde{G}}-F}\right)^{ab}\,,$ $\Theta^{ab}=\left(\frac{1}{{\tilde{G}}+F}F\frac{1}{{\tilde{G}}-F}\right)^{ab}\,.$
The symmetric part is the inverse of the following metric:
@@
On the other hand, if one chooses another scheme for the basis (19), it is possible to change the coefficient of the last term in the second line to be 2, and remove the first term in the fifth line of (32). Then all terms in (32) with the structure $\Omega_{ab}{}^{\mu}\Omega_{cd\mu}$ whose world-volume indices are contracted with an even number of $F$, i.e., all terms in the first and second lines and the first term in the third line, can be written as $-{\tilde{{\cal G}}}^{ac}{\tilde{{\cal G}}}^{bd}\Omega_{ab}{}^{\mu}\Omega_{cd\mu}$. Therefore, one may expect that the world-volume covariant derivative constructed from the above metric, i.e., $\nabla_{{\tilde{{\cal G}}}}$, may be the appropriate covariant derivative to consider for constructing the second fundamental form and constructing the independent basis in which all indices are contracted with ${\tilde{{\cal G}}}^{ab}$ and $\Theta^{ab}$. Then T-duality can fix the coupling constants of such a ${\tilde{{\cal G}}}$-covariant basis. It would be interesting to perform the appropriate calculation to determine the coupling constants of the ${\tilde{{\cal G}}}$-covariant basis. It may be possible that the infinite number of couplings at order $\alpha^{\prime}$ can be expressed as a few couplings constructed from the ${\tilde{{\cal G}}}$-metric and $\Theta^{ab}$.